OpenStudy (anonymous):

if cosθ=1/4, find tanθ?

OpenStudy (anonymous):

In which quadrants is cosine positive?

OpenStudy (anonymous):

Are there any constraints on tangent?

OpenStudy (anonymous):

cosine is positive in quadrants 1 and 4

OpenStudy (anonymous):

and θ is between 0 and 90 degrees

OpenStudy (anonymous):

\[find (\sin \theta) first ---->\sin \theta=(\sqrt{1-\cos^2 \theta})\]

OpenStudy (anonymous):

now \[\tan \theta=\sin \theta/\cos \theta\]

OpenStudy (anonymous):

tan theta is positive cuz 0-90 degrees means first quadrant..